OpenStudy (anonymous):

use langrange multipliers to find the extrema of f subject to the stated constraints. f(xyz) = x + y + z constraint = x^2+y^2+z^2=25

OpenStudy (anonymous):

this is what i've gotten so far and i'm not sure if i'm doing it right / where to go from here... \[x+y+z - \lambda (x^2+y^2+z^2-25) \] \[x+y+z-\lambda x^2-\lambda y^2-\lambda z^2 +25\lambda\] \[Fx = 1-2x \lambda \] \[fy= 1 - 2y \lambda\] \[fz= 1 - 2z \lambda\] \[f \lambda -x^2 - y^2-z^2 +25\] \[x = 1/2\lambda\] \[y = 1/2\lambda\] \[z=1/2\lambda\] \[\lambda = 1/33.3\]

OpenStudy (anonymous):

please halp ._.

OpenStudy (experimentx):

\[ \nabla f(x,y,z) = \lambda \nabla(x^2 + y^2 + z^2)\]

OpenStudy (anonymous):

yeah, that's what i did to get x+y+z−λx2−λy2−λz2+25λ

OpenStudy (experimentx):

\[ <1, 1, 1> = \lambda <2x, 2y,2z> \]

OpenStudy (anonymous):

wait.. why is it < 1 1 1 > ??

OpenStudy (anonymous):

nvm, its the derivates.. you're taking the derivatives first

OpenStudy (experimentx):

\( \nabla \) is gradient operator

OpenStudy (anonymous):

my book / professor has us do the equation out first, and then separately find fx, fy, fz, and flambda is there a reason why you've done it backwards? does this make it easier? how do i find the extrema with this?

OpenStudy (experimentx):

I am not so sure ... i've just learned a day back. http://ocw.mit.edu/courses/mathematics/18-02-multivariable-calculus-fall-2007/video-lectures/lecture-13-lagrange-multipliers/ If you try to visualise at maximum or minimum, the normal to curves are parallel to each other .... it makes much sense. They say so .... and i agree!!

OpenStudy (experimentx):

In this case, it is minimization of function!!

OpenStudy (anonymous):

how do you know that? :o and how do i do that?

OpenStudy (experimentx):

picture ... of the function and constraint. a plane and a sphere (i guess both maximum and minimum value are same ... in this case ... because of symmetricity)

OpenStudy (anonymous):

i'm not sure i follow but.. the answer is min f(-5/root3, -5/root3, -5/root3) = -5/root3 max: same but everything is positive.. is my value of lambda incorrect? =\

OpenStudy (anonymous):

nvms i gots it ^.^ thanks!